# Knowledge Boundary of Large Language Models: A Survey

### v1

@@ 5.2 Model-specific Unknown Knowledge
5.2.1 External Knowledge Retrieval
@@
5.2.2 Parametric Knowledge Editing
Researchers also develop knowledge editing methods for altering model behaviors to modify specific parameters within the LLM without extensive retraining. According to the memory mechanism, we categorize existing knowledge editing methods into three categories: explicit memory space, implicit memory space, and no memory space. As for explicit memory space, these approaches Mitchell et al. (2022); Zheng et al. (2023a); Madaan et al. (2022); Song et al. (2024b); Zhong et al. (2023) use a memory pool to retrieve and apply edits via prompts. As for implicit memory space, these approaches activate the LLM’s parametric memory space based on specific input triggers, such as codebook Hartvigsen et al. (2023), neurons Huang et al. (2023c); Dong et al. (2022), LoRA blocks Yu et al. (2024), and FFN side memories Wang et al. (2024c). Another group of methods does not adopt extra memory components. Instead, they adopt various techniques to directly edit the original model parameters, such as meta learning Tan et al. (2024) and locate-then-edit Meng et al. (2022, 2023).
5.2.3 Knowledge-enhanced Fine-tuning
Knowledge-enhanced fine-tuning internalizes new knowledge into models by leveraging structured or synthetic representations. This involves encoding knowledge as factual records, synthetic corpora, and domain-specific taxonomies. Techniques such as fact-based encoding Mecklenburg et al. (2024), synthetic data creation Joshi et al. (2024), and hierarchical organization Liu et al. (2024c) ensure comprehensive domain coverage, while interleaved generation and context-aware structuring Zhang et al. (2024b) aim to enhance the data quality.
@@ 5.3.1 Refusal
Prompt-based Approaches
use designed prompts that help LLMs decide whether to refuse questions about unknown knowledge. The prompts are used to evaluate if a question involves unknown content to LLM Wen et al. (2024a); Amayuelas et al. (2024); Agarwal et al. (2023), and to express the knowledge limitations Chen et al. (2024c). Also, LLMs can be prompted to justify their decision to decline a question Song et al. (2024a).
Alignment-based Approaches
include supervised fine-tuning and reinforcement learning approaches. Supervised methods involve creating honesty alignment datasets, such as “I don’t know” datasets, to teach LLMs to admit uncertainty in responses Yang et al. (2023b); Cheng et al. (2024b); Zhang et al. (2024a). Reinforcement learning approaches generally constructs datasets that reflect user preferences, and use them to train LLMs through reward systems to discern when to refuse questions Cheng et al. (2024b); Tomani et al. (2024); Xu et al. (2024a).
5.3.2 Asking Clarification Questions
When LLMs encounter questions involving model-agnostic unknown knowledge, asking clarification questions is an another common strategy. This method avoids direct uncertain responses and uses proactive dialogues to refine queries Deng et al. (2023a); Aliannejadi et al. (2021); Guo et al. (2021); Leippert et al. (2024). This is supported by specific prompt frameworks, with schemes encouraging LLMs to analyze questions deeply before responding Deng et al. (2023b); Chen et al. (2024f). Frameworks by Kuhn et al. (2022) and Mu et al. (2023) enable LLMs to request clarifications selectively or identify unclear requirements, enhancing response accuracy. Latest methods like contrastive self-training and reward model learning help improve the quality of LLMs’ questions in dialogues Chen et al. (2024d); Andukuri et al. (2024).
@@ 6 Challenges and Prospects
Various knowledge-based QA datasets are key benchmarks for assessing LLMs’ knowledge boundaries, as summarized in Appendix B. However, determining ground truth is challenging, as failures may stem from a lack of knowledge, poor prompts, or complicated reasoning. Additionally, failing to answer a single question does not necessarily indicate whether the LLM can handle related knowledge Yin et al. (2024). There is a pressing need for more comprehensive benchmarks to effectively evaluate the knowledge boundaries of LLMs.
@@
While knowledge boundary studies are often conducted in specific domains, understanding the general knowledge boundary in LLMs is vital. The internal state probing approach has been validated with a certain generalization ability Liu et al. (2024b), but it is still an open challenge whether trained probes can generalize well across domains as a general knowledge boundary detector, fostering refusal and input clarification in open domains. Further theoretical analysis and studies are needed to identify the existence and utility of general knowledge boundaries.
Utilization of Knowledge Boundary
Estimating and understanding LLMs’ knowledge boundaries should not mark the end of the process. Instead, identifying these limitations can serve as a foundation for enhancing the model’s performance in mitigating queries beyond their knowledge boundaries. For instance, the utilization of model uncertainty can reduce RAG costs and minimize the risk of introducing noise from external sources Yao et al. (2024), or enhance the preference optimization by encouraging the LLM policy to differentiate reliable or unreliable feedback Wang et al. (2024a).
@@ 7 Conclusions
This survey presented a comprehensive overview of the knowledge boundary of LLMs, offering a formalized taxonomy and addressing key challenges in the field. By exploring undesirable behaviors, identification techniques, and mitigation strategies, we emphasized the critical role of understanding and managing these boundaries to improve the reliability and utility of LLMs. Despite significant progress, challenges persist, including lack of comprehensive benchmarks, achieving domain generalization, potential uses of knowledge boundaries, and addressing unintended side effects. We hope this survey inspires continued exploration and innovation toward more trustworthy and reliable LLMs.
@@
This survey does not give a formal definition of the knowledge $k$, which is a critical problem in the scope of NLP research on knowledge. In this survey, we define the abstracted concept of knowledge as $k$, which is represented by a set of textual expressions of input and output. This definition can facilitate practical NLP experiments and efficient validation. In fact, the formal definition of knowledge is still a debatable topic, calling for future exploration. For example, Fierro et al. (2024) try to bridge the philosophical definition to the knowledge of LLMs, though significant disagreements persist among various philosophical schools of thought.
@@
Besides, in our definition, the known knowledge for LLM lies within the universal knowledge boundary, which is the knowledge known for human. We generally believe that LLMs do not possess knowledge beyond this boundary. However, there may be outliers that LLMs have knowledge that is unknown for human, which is not clearly studied in existing research. Wang et al. (2024b) hypothesize that LLMs may create new knowledge, but the creation may be unreliable, remaining an open question.
@@ Appendix A Overview
Controversial Knowledge , color=cyan!100, fill=cyan!40, text=black [ Wang et al. (2024e); Amayuelas et al. (2024); Singh et al. (2024); Naous et al. (2024) , behaviour_leaf, text width=34.2em ] ] ] ] [ Identification
@@
Decomposition, color=lightgreen!100, fill=lightgreen!40, text=black [ Hou et al. (2024); Ling et al. (2024); Yadkori et al. (2024); Gao et al. (2024); Aichberger et al. (2024),
Ling et al. (2024); Ahdritz et al. (2024) , identification_leaf, text width=34.2em ] ] [ Token Probability-based
Uncertainty Estimation, color=lightgreen!100, fill=lightgreen!40, text=black [ Manakul et al. (2023); Huang et al. (2023b); Duan et al. (2023); Fadeeva et al. (2024) , identification_leaf, text width=34.2em ] ] [ Semantic-based
Uncertainty Estimation, color=lightgreen!100, fill=lightgreen!40, text=black [ Kuhn et al. (2023); Lin et al. (2024); Chen and Mueller (2024); Manakul et al. (2023); Zhao et al. (2024c),
@@
Calibration, color=lightgreen!100, fill=lightgreen!40, text=black [ Tao et al. (2024); Liu et al. (2024d); Xie et al. (2024); Shen et al. (2024); Mielke et al. (2022); Stengel-Eskin et al. (2024) , identification_leaf, text width=34.2em ] ] ] [ Internal State Probing, color=lightgreen!100, fill=lightgreen!60, text=black [ Burns et al. (2023); Kossen et al. (2024); Li et al. (2024a); Azaria and Mitchell (2023); Ji et al. (2024); Liu et al. (2024b); Marks and Tegmark (2023) , identification_leaf, text width=42.5em ] ] ] [ Mitigation, color=lightpurple!100, fill=lightpurple!100, text=black [ Prompt-sensitive
Known Knowledge, color=lightpurple!100, fill=lightpurple!60, text=black [ Prompt Optimization, color=lightpurple!100, fill=lightpurple!30, text=black [ Zhou et al. (2023b); Li et al. (2023a); Yang et al. (2024c); Wang et al. (2024d); Prasad et al. (2023),
@@
Reasoning, color=lightpurple!100, fill=lightpurple!30, text=black [ Wei et al. (2022); Zhou et al. (2023a); Yao et al. (2023); Zheng et al. (2023b); Press et al. (2023); Wang et al. (2022),
Jung et al. (2022); Liu et al. (2022b); Zhang et al. (2024c); Wu et al. (2024a); Zhao et al. (2023); Li et al. (2024e),
Trivedi et al. (2023) , mitigation_leaf, text width=34.2em ] ] [ Self-refinement, color=lightpurple!100, fill=lightpurple!30, text=black [ Madaan et al. (2024); Miao et al. (2024); Manakul et al. (2023); Weng et al. (2023); Dhuliawala et al. (2024),
Huang et al. (2024); Li et al. (2024b); Du et al. (2024); Fu et al. (2023) , mitigation_leaf, text width=34.2em ] ] [ Factuality Decoding, color=lightpurple!100, fill=lightpurple!30, text=black [ Bi et al. (2024); Li et al. (2023b); Chuang et al. (2024); Chen et al. (2024b); Kai et al. (2024); Yang et al. (2024b),
@@
Hartvigsen et al. (2023); Huang et al. (2023c); Dong et al. (2022); Yu et al. (2024); Wang et al. (2024c),
@@
Fine-tuning, color=lightpurple!100, fill=lightpurple!30, text=black [ Mecklenburg et al. (2024); Joshi et al. (2024); Liu et al. (2024c); Zhang et al. (2024b), , mitigation_leaf, text width=34.2em ] ] ] [ Model-agnostic
@@
Cheng et al. (2024b); Zhang et al. (2024a); Cheng et al. (2024b); Tomani et al. (2024); Xu et al. (2024a) , mitigation_leaf, text width=34.2em ] ] [ Asking Clarification

### v2

@@ 5.2 Model-specific Unknown Knowledge
External Knowledge Retrieval
@@
Parametric Knowledge Editing
Researchers also develop knowledge editing methods for altering model behaviors to modify specific parameters within the LLM without extensive retraining. According to the memory mechanism, we categorize existing knowledge editing methods into three categories: explicit memory space, implicit memory space, and no memory space. As for explicit memory space, these approaches Mitchell et al. (2022); Zheng et al. (2023a); Madaan et al. (2022); Song et al. (2024b); Zhong et al. (2023) use a memory pool to retrieve and apply edits via prompts. As for implicit memory space, these approaches activate the LLM’s parametric memory space based on specific input triggers, such as codebook Hartvigsen et al. (2023), neurons Huang et al. (2023d); Dong et al. (2022), LoRA blocks Yu et al. (2024), and FFN side memories Wang et al. (2024d). Another group of methods does not adopt extra memory components. Instead, they adopt various techniques to directly edit the original model parameters, such as meta learning Tan et al. (2024) and locate-then-edit Meng et al. (2022, 2023).
Knowledge-enhanced Fine-tuning
Knowledge-enhanced fine-tuning internalizes new knowledge into models by leveraging structured or synthetic representations. This involves encoding knowledge as factual records, synthetic corpora, and domain-specific taxonomies. Techniques such as fact-based encoding Mecklenburg et al. (2024), synthetic data creation Joshi et al. (2024), and hierarchical organization Liu et al. (2025) ensure comprehensive domain coverage, while interleaved generation and context-aware structuring Zhang et al. (2024b) aim to enhance the data quality.
@@ 5.3 Model-agnostic Unknown Knowledge
Prompt-based approaches use designed prompts that help LLMs decide whether to refuse questions about unknown knowledge. The prompts are used to evaluate if a question involves unknown content to LLM Wen et al. (2024a); Amayuelas et al. (2024); Agarwal et al. (2023), and to express the knowledge limitations Chen et al. (2024c). Also, LLMs can be prompted to justify their decision to decline a question Song et al. (2024a).
Alignment-based approaches include supervised fine-tuning and reinforcement learning (RL) approaches. Supervised methods involve creating honesty alignment datasets, such as “I don’t know” datasets, through instruct tuning to teach LLMs to admit uncertainty in responses Yang et al. (2023b); Cheng et al. (2024b); Zhang et al. (2024a); Gao et al. (2024a); Zhu et al. (2025). RL approaches generally constructs datasets that reflect user preferences, and use them to train LLMs through reward systems to discern when to refuse questions Cheng et al. (2024b); Tomani et al. (2024); Xu et al. (2024a).
Asking Clarification Questions
When LLMs encounter questions involving model-agnostic unknown knowledge, asking clarification questions is an another common strategy. This method avoids direct uncertain responses and uses proactive dialogues to refine queries Deng et al. (2023a); Aliannejadi et al. (2021); Guo et al. (2021); Leippert et al. (2024); Kim et al. (2024a). This is supported by specific prompt frameworks, with schemes encouraging LLMs to analyze questions deeply before responding Deng et al. (2023b); Chen et al. (2024f). Frameworks by Kuhn et al. (2022) and Mu et al. (2023) enable LLMs to request clarifications selectively or identify unclear requirements, enhancing response accuracy. Latest methods like contrastive self-training and reward model learning help improve the quality of LLMs’ questions in dialogues Chen et al. (2024d); Andukuri et al. (2024).
@@ 6 Challenges and Prospects
Various knowledge-based QA datasets are key benchmarks for assessing LLMs’ knowledge boundaries, as summarized in Appendix B. However, there are still critical areas lacking comprehensive benchmarks. Firstly, it lacks benchmarks for identifying the knowledge boundary of LLMs (§4). The benchmark construction should involve key aspects including multiple ground-truth answers, the influence of prompts, and reasoning complexity. Failing to answer a single question does not necessarily indicate whether the LLM can handle related knowledge Yin et al. (2024). Secondly, evaluating mitigation methods under different categories (§5) also requires corresponding benchmarks. A standardized benchmark is essential for enabling a thorough and fair comparison on the performance of various mitigation methods. Thus, our proposed taxonomy provides a systematic and valuable foundation to guide the development of these benchmarks.
Utilization of Knowledge Boundary
Estimating and understanding LLMs’ knowledge boundaries should not mark the end of the process. Instead, identifying these limitations can serve as a foundation for enhancing the model’s performance in mitigating queries beyond their knowledge boundaries. For instance, the utilization of model uncertainty can reduce RAG costs and minimize the risk of introducing noise from external sources Yao et al. (2024), or enhance the preference optimization by encouraging the LLM policy to differentiate reliable or unreliable feedback Wang et al. (2024a). Another instance is to enhance the robustness of LLMs against prompt sensitivity. Some pioneer research study such issue regarding the order of demonstrations CHEN et al. (2025); Lu et al. (2022). Further studies could investigate the role of outward knowledge boundary of LLMs in the overall prompt robustness of LLMs, enabling them to express more knowledge they already possess.
Understanding Knowledge Boundary through Knowledge Mechanisms
Existing research on knowledge mechanisms, including memorization, comprehension, creation, and evolution, investigates how LLMs acquire, store, and utilize knowledge Wang et al. (2024c). It is worth studying different phenomena of LLM knowledge boundaries under these mechanism views. For example, the outward knowledge boundary showcases how mechanisms like memorization and comprehension manifest in the explicit behaviors and outputs of the model, while the parametric boundary reflects a deeper, less visible level of how knowledge is embedded and structured through these mechanisms. The universal boundary can help measure the creative and evolutionary capabilities of LLMs.
@@
While knowledge boundary studies are often conducted in specific domains, understanding the general knowledge boundary in LLMs is vital. The internal state probing approach has been validated with a certain generalization ability Liu et al. (2024b), but it is still an open challenge whether trained probes can generalize well across domains as a general knowledge boundary detector, fostering refusal and input clarification in open domains. Further theoretical analysis and studies are needed to identify the existence and utility of general knowledge boundaries, which may be related to fundamental theories of LLM knowledge mechanism Wang et al. (2024b); Allen-Zhu and Li (2024).
@@
Knowledge Boundary in Long-Form Language Modeling
Knowledge boundaries critically impact long-form factuality, defining how well LLMs generate coherent and accurate extended responses. Unlike short-form factuality, which depends on individual fact retrieval, long-form factuality is affected by cumulative knowledge gaps, where minor errors propagate over extended discourse. Existing research (Wei et al., 2024; Min et al., 2023; Huang and Chen, 2024a) has explored evaluation and mitigation strategies, providing a lens to examine how LLMs navigate and extend their knowledge boundaries, but the interaction between knowledge boundaries and factuality degradation remains an open research area.
@@ 7 Conclusions
This survey present a comprehensive overview of the knowledge boundary of LLMs, offering a formalized taxonomy and addressing key questions in the field. By exploring undesirable behaviors, identification techniques, and mitigation strategies, we emphasize the critical role of understanding and managing these boundaries to improve the reliability and utility of LLMs. Despite significant progress, challenges persist, including lack of comprehensive benchmarks, potential uses of knowledge boundary, and the role of knowledge boundary under various mechanisms. We hope this survey inspires continued exploration and innovation toward more trustworthy and reliable LLMs.
@@
This survey does not give a formal definition of the knowledge $k$, which is a critical problem in the scope of NLP research on knowledge. In this survey, we define the abstracted concept of knowledge as $k$, which is represented by a set of textual expressions of input and output $Q_{k}$. This definition can facilitate practical NLP experiments and efficient validation. In fact, the formal definition of knowledge is still a debatable topic, calling for future exploration. For example, Fierro et al. (2024) try to bridge the philosophical definition to the knowledge of LLMs, though significant disagreements persist among various philosophical schools of thought.
Various Forms of Textual Expressions regarding Different Knowledge Types
Different types of knowledge may correspond to various forms of textual input-output $Q_{k}$, while we aim to provide a universal definition without the loss of generality. For instance, outputs for complex concepts may be open-ended and long-form, while simpler concepts might be expressed in a multiple-choice format. Some knowledge can be explicitly stated in the input, whereas others, such as commonsense knowledge, may need to be inferred from the input. Additionally, a single input may have multiple valid outputs. Some knowledge types, like mathematical knowledge, may inevitably involve multiple pieces of knowledge in a single input-output instance. As research progresses, a more nuanced definition for $Q_{k}$ may be necessary to accommodate different knowledge types effectively.
@@
Besides, in our definition, LLMs operate within the universal knowledge boundary, typically limited to human-known knowledge. We generally believe that LLMs do not possess knowledge beyond this boundary. However, there may be outliers that LLMs have knowledge that is unknown for human, which is not clearly studied in existing research. Wang et al. (2024b) hypothesize that LLMs may create new knowledge, but its reliability remains uncertain. While such outputs could reflect meaningful discoveries, they may also stem from implicit correlations in training data. Since existing research has not systematically examined LLM-generated unknown knowledge, its nature and implications remain unclear.
Acknowledgment
This research was supported by the Singapore Ministry of Education (MOE) Academic Research Fund (AcRF) Tier 1 grant (No. MSS24C012).
@@ Appendix A Overview
Controversial Knowledge , color=cyan!100, fill=cyan!40, text=black [ Wang et al. (2024f); Amayuelas et al. (2024); Singh et al. (2024); Naous et al. (2024) , behaviour_leaf, text width=34.2em ] ] ] ] [ Identification
@@
Decomposition, color=lightgreen!100, fill=lightgreen!40, text=black [ Hou et al. (2024); Ling et al. (2024); Yadkori et al. (2024); Gao et al. (2024b); Aichberger et al. (2024),
Ling et al. (2024); Ahdritz et al. (2024) , identification_leaf, text width=34.2em ] ] [ Conformal
Predicion, color=lightgreen!100, fill=lightgreen!40, text=black [ Kumar et al. (2023); Cherian et al. (2024); Mohri and Hashimoto (2024); Su et al. (2024); Ravfogel et al. (2023); Ye et al. (2024),
Zollo et al. (2024); Quach et al. (2024); Li et al. (2024d) , identification_leaf, text width=34.2em ] ] [ Token Probability-based
Uncertainty Estimation, color=lightgreen!100, fill=lightgreen!40, text=black [ Manakul et al. (2023); Huang et al. (2023c); Duan et al. (2023); Fadeeva et al. (2024) , identification_leaf, text width=34.2em ] ] [ Semantic-based
Uncertainty Estimation, color=lightgreen!100, fill=lightgreen!40, text=black [ Kuhn et al. (2023); Lin et al. (2024b); Chen and Mueller (2024); Manakul et al. (2023); Zhao et al. (2024c),
@@
Calibration, color=lightgreen!100, fill=lightgreen!40, text=black [ Tao et al. (2024); Liu et al. (2024c); Xie et al. (2024); Shen et al. (2024); Mielke et al. (2022); Stengel-Eskin et al. (2024) , identification_leaf, text width=34.2em ] ] ] [ Internal State Probing, color=lightgreen!100, fill=lightgreen!60, text=black [ Burns et al. (2023); Kossen et al. (2024); Li et al. (2024a); Azaria and Mitchell (2023); Ji et al. (2024); Liu et al. (2024b); Marks and Tegmark (2023) , identification_leaf, text width=42.5em ] ] ] [ Mitigation, color=lightpurple!100, fill=lightpurple!100, text=black [ Prompt-sensitive
Known Knowledge, color=lightpurple!100, fill=lightpurple!60, text=black [ Prompt Optimization, color=lightpurple!100, fill=lightpurple!30, text=black [ Zhou et al. (2023b); Li et al. (2023b); Yang et al. (2024c); Wang et al. (2024e); Prasad et al. (2023),
@@
Reasoning, color=lightpurple!100, fill=lightpurple!30, text=black [ Wei et al. (2022b); Zhou et al. (2023a); Yao et al. (2023); Zheng et al. (2023b); Press et al. (2023); Wang et al. (2022),
Jung et al. (2022); Liu et al. (2022b); Zhang et al. (2024c); Wu et al. (2024a); Zhao et al. (2023); Li et al. (2024f) , mitigation_leaf, text width=34.2em ] ] [ Self-refinement, color=lightpurple!100, fill=lightpurple!30, text=black [ Madaan et al. (2024); Miao et al. (2024); Manakul et al. (2023); Weng et al. (2023),
Huang et al. (2024b); Li et al. (2024b); Du et al. (2024); Fu et al. (2023) , mitigation_leaf, text width=34.2em ] ] [ Factuality Decoding, color=lightpurple!100, fill=lightpurple!30, text=black [ Bi et al. (2024); Li et al. (2023c); Chuang et al. (2024); Chen et al. (2024b); Kai et al. (2024); Yang et al. (2024b),
@@
Hartvigsen et al. (2023); Huang et al. (2023d); Dong et al. (2022); Yu et al. (2024); Wang et al. (2024d),
@@
Fine-tuning, color=lightpurple!100, fill=lightpurple!30, text=black [ Mecklenburg et al. (2024); Joshi et al. (2024); Liu et al. (2025); Zhang et al. (2024b), , mitigation_leaf, text width=34.2em ] ] ] [ Model-agnostic
@@
Cheng et al. (2024b); Zhang et al. (2024a); Zhu et al. (2025); Tomani et al. (2024); Xu et al. (2024a) , mitigation_leaf, text width=34.2em ] ] [ Asking Clarification
@@
Summary of Contribution
As a survey paper, our primary goal is to synthesize and analyze existing research while providing new insights and frameworks for understanding the knowledge boundaries of LLMs. We believe our work offers significant novelty in the following aspects:
1) Scope and Coverage
  • Novelty in Scope: This survey covers a topic or area that has not been thoroughly reviewed before. We address an emerging field and underexplored topics.
  • Comprehensiveness: This survey provides a comprehensive or up-to-date overview. Novelty lies in including recent advancements, overlooked studies, or a broader range of perspectives.
2) Organization and Structure
  • Unique Frameworks or Taxonomies: This survey introduces a novel taxonomy, as a new way of categorizing, organizing, or analyzing the literature of LLM knowledge boundary studies. This new taxonomy provides fresh insights into the field.
3) Insights and Critical Analysis
  • Original Insights: This survey provides original interpretations and thought-provoking perspectives on the existing literature. For instance, we provide a unique categorization of mitigation strategies for prompt-sensitive known knowledge according to the process on the LLM input and output sides, and also for model-specific unknown knowledge based on the extent of modification to LLM’s parameters.
  • Identification of Gaps: This survey identifies several underexplored areas or open problems in the field.
4) Timeliness and Relevance
  • As the very first survey paper on the topic of LLM knowledge boundaries, our work serves as a foundational resource for researchers and practitioners. By consolidating and organizing the existing literature, we provide a starting point for further exploration and innovation in this critical area.
In summary, our work contributes novelty through comprehensive coverage, innovative taxonomy, and original insights. We believe these contributions significantly advance the understanding of LLM knowledge boundaries and provide a valuable resource for the research community.